### v1

# IPR-1: Interactive Physical Reasoner

Abstract
Humans learn by observing, interacting with environments, and internalizing physics and causality. Here, we aim to ask whether an agent can similarly acquire human-like reasoning from interaction and keep improving with more experience. We study this in a Game-to-Unseen (G2U) setting, curating 1,000+ heterogeneous games with diverse physical and causal mechanisms, and evaluate at three human-like levels: Survival, Curiosity, Utility, from primitive intuition to goal-driven reasoning. Our analysis reveals complementary failures: VLM/VLA agents reason but lack look-ahead in interactive settings, while world models imagine but imitate visual patterns rather than analyze physics and causality. We therefore propose IPR (Interactive Physical Reasoner), using world-model rollouts to score and reinforce a VLM’s policy, and introduce PhysCode, a physics-centric action code aligning semantic intent with dynamics to provide a shared action space for prediction and reasoning. Pretrained on 1,000+ games, our IPR performs robustly on three levels, matches GPT-5 overall, and surpasses it on Curiosity. We find that performance improves with more training games and interaction steps, and that the model also zero-shot transfers to unseen games. These results support physics-centric interaction as a path to steadily improving physical reasoning. Our code will be publicly available.
footnotetext: Equal contribution.  Corresponding author. ∗∗Work conducted at Shanghai Jiao Tong University.
1 Introduction
Humans do not learn physics and causality from labels; we earn them through interaction. As experience accumulates with age, our prediction sharpens, our reasoning stabilizes, and our abilities scale. This motivates a central question for embodied AI: what learning paradigm enables human-like reasoning to learn through interactive experience, and to improve steadily with more interaction?
We assume that, if an agent is exposed to diverse, interactive worlds and trained to capture shared physical and causal mechanisms, rather than domain-specific appearance or action interfaces, it would scale its physical reasoning ability reliably and transfer to new scenarios. This view aligns with prior reasoning works [Bengio+chapter2007, huang2023reasoninglargelanguagemodels, zhang2024stepbackrethinkingstages]. Large VLM/VLA models [wang2024jarvis, wang2025gametarspretrainedfoundationmodels, bu2025univlalearningacttaskcentric], pretrained on vast corpora, exhibit basic reasoning abilities and a certain degree of transfer; yet in interactive scenarios they lack forward prediction visually and often fail when precise anticipation of action consequences is required (e.g., timing, contact, momentum) (Fig. 2(b)). World models and other prediction-based methods [bruce2024genie, assran2025v, hafner2025trainingagentsinsidescalable] can imagine futures via differentiable latent dynamics and interactively optimize trajectories toward goal-aligned representations, but they tend to collapse into target-chasing imitation rather than genuine causal reasoning, failing in long-horizon and sparse-reward tasks. This naturally raises a question: can we blend the open-ended reasoning of VLMs with the predictive grounding of world models to support interactive reasoning in novel environments and yield competence that improves steadily with experience?
Figure 1: Game-to-Unseen (G2U) problem. Humans accumulate interactive experience and rapidly adapt to new games. Despite different visuals and interfaces, many games share underlying physical/causal mechanisms. We pretrain on 1,000+ visually and physically diverse games to test whether an agent can internalize these shared mechanisms and generalize to unseen games.
(a) Three-level evaluation inspired by Maslow’s hierarchy of needs. We organize tasks into a pyramid of Survival, Curiosity, and Utility. Survival measures how long the agent can stay alive by avoiding risks; Curiosity measures how broadly it visits novel states; and Utility measures how well it achieves downstream goals. The three levels progress from physical intuition to goal-driven reasoning. Our IPR performs robustly across the entire pyramid.
(b) Motivating failure cases in control semantics, language grounding, and prediction. (1) Control conflict: the same key (e.g., UP) triggers different semantics across games (camera tilt up v.s. character move up), causing console aliasing. (2) Vision-language distortion: text-only actions cannot specify precise visual magnitudes (e.g., jump height/speed), leading to systematic amplitude errors. (3) Missing foresight: without imagination, the agent cannot anticipate upcoming hazards during interaction (e.g., spikes, moving enemies).
Figure 2: Overview: three-level evaluation pyramid (left) and failure cases of previous VLM-based model (right), motivating our IPR.
In this way, we propose IPR (Interactive Physical Reasoner): a paradigm where world model prediction reinforces a VLM policy to adapt its physical reasoning in interactive environments (Fig. 3). A natural obstacle arises when naively piping VLM outputs into a predictor: keyboard induces interface mismatches across games and language distorts visual details (Fig. 2(b)). We therefore introduce PhysCode, a physics-centric action code that fuses action semantics with visual dynamics into a compact discrete representation. Concretely, we follow a Genie [bruce2024genie] style discretization over fused features extracted from video frames, optical flow, and action semantics. Each code is encouraged to align with (i) domain-agnostic dynamical primitives (e.g., momentum change) and (ii) domain-specific visual affordances. Instead of issuing raw keys or free-form language, the reasoning policy outputs multiple PhysCode sequences, which are scored by the world model in the same latent space so that the best candidate would be executed and its imagined rewards are used to reinforce the policy.
To evaluate the paradigm at scale, we curate over 1,000 heterogeneous games spanning visual styles, control interfaces, and physical and causal mechanisms. Games form a low-cost, controllable testbed for physical reasoning: they afford rich interaction, physics closely resembling the real world, and effectively unlimited rollouts. We further organize evaluation into three levels inspired by Maslow’s hierarchy of needs [huitt2007maslow]: Survival, Curiosity, Utility, covering a spectrum from physical intuition to goal-directed reasoning (Fig. 2(a)). The result on three levels verifies two failure modes: reasoning-based VLM/VLA lack forward consequence prediction to explore (Curiosity), while prediction-based world models explore broadly yet fail at goal-driven tasks (Utility), which motivates our design.
Across this suite, our IPR remains robust on all three levels, while RL-based and prediction-based baselines often collapse on one or more of them. Trained under the IPR paradigm, an 8B backbone matches GPT-5 overall and even surpasses it on curiosity. Moreover, competence scales with the number of training games and interaction steps (Fig. 5) and zero-shot transfers to novel environments, highlighting the potential of interactive learning for physical reasoning at scale.
In general, our contributions are: (1) We formulate the G2U problem and curate 1,000+ heterogeneous games with a hierarchical evaluation (Survival/Curiosity/Utility), diagnosing the strengths and weaknesses of prevalent prediction-based, RL-based, and VLM-based methods. (2) We propose IPR: world-model rollouts score and reinforce VLM in the same action space, enabling interactive experience to steadily build up physical reasoning ability. (3) We introduce PhysCode, a physics-centric action code fusing action semantics with visual dynamics, bridging WM prediction and VLM reasoning.
2 Related Works
Action Space Discovery.
Research on action spaces spans hand-designed controls, language-based interfaces, and learned latent representations. Early embodied agents operated over environment-specific key bindings, torques, or joystick signals [mnih2015human, schulman2017proximal, haarnoja2018softactorcriticoffpolicymaximum, brockman2016openaigym], which offer precise control but entangle behavior with platform-specific layouts and hinder cross-domain transfer. A second line adopts language-based action spaces, issuing natural-language commands or tool calls [shridhar2020alfredbenchmarkinterpretinggrounded, fan2022minedojobuildingopenendedembodied, wang2024jarvis, wang2023voyager, ahn2022icanisay]; while language affords semantic generality, it abstracts away timing, force, and perception–action couplings, often leading to imprecise or under-grounded control [savva2019habitat, reed2022generalist]. A complementary direction learns latent action spaces directly from interaction data. Discrete or continuous latent codes—via VQ-VAE [oord2018neuraldiscreterepresentationlearning] or sequence models—have been explored for planning, control, and world models [bruce2024genie, chi2024diffusionpolicyvisuomotorpolicy, sharma2020emergentrealworldroboticskills, lynch2021languageconditionedimitationlearning, okada2021dreamingmodelbasedreinforcementlearning]. Recent VLM/VLA systems integrate such latent tokens into large multimodal models [sapkota2025vision, huang2024embodiedgeneralistagent3d], but these codes often remain entangled across domains and lack mechanisms to capture shared physical principles versus environment-specific affordances. Our work addresses this gap by learning a physics-centric latent action space that captures reusable dynamical patterns across games, instead of binding actions to domain-specific visuals and control layouts.
Agents in Interactive Environments.
Research on game-playing agents has largely followed three threads. RL-based agents, from DQN and PPO/SAC to large-scale systems like AlphaStar and OpenAI Five [mnih2013playing, vanhasselt2015deepreinforcementlearningdouble, schulman2017proximal, haarnoja2018softactorcriticoffpolicymaximum, Vinyals2019GrandmasterLI, openai2019dota2largescale], learn policies directly from pixels and rewards and achieve strong title-specific performance, but remain sample-inefficient, brittle to interface changes, and struggle with long-horizon credit assignment and cross-game transfer. Prediction-based (world-model) agents such as World Models, PlaNet, the Dreamer family, and Genie [https://doi.org/10.5281/zenodo.1207631, hafner2019learninglatentdynamicsplanning, hafner2020dreamcontrollearningbehaviors, hafner2023mastering, bruce2024genie] first learn latent dynamics and then plan or optimize in imagination, improving exploration and sparse-reward learning, yet degrade when learned dynamics or action semantics drift from the test environment and typically optimize task or pixel losses rather than reasoning quality. VLM/VLA-based agents like Gato, RT-2, Voyager, MineDojo, and recent VLA frameworks [reed2022generalist, brohan2023rt2visionlanguageactionmodelstransfer, wang2023voyager, fan2022minedojobuildingopenendedembodied] cast acting as sequence modeling over images, text, and actions and excel at zero-shot instruction following, but rely heavily on static corpora, heuristic wrappers, and weakly grounded forward prediction. Our IPR paradigm aims to inherit the strengths of these lines by using a physics-centric latent action space where a world model provides imagination-based value estimates and a reasoning VLM policy is reinforced through interactive experience in the same latent space.
Benchmarks and Evaluation.
Interactive environments have long served as testbeds for learning control, exploration, and generalization: Atari/ALE provided dense stepwise rewards for RL training and evaluation [mnih2013playing, bellemare2013arcade], while later platforms such as Minecraft, VizDoom, and StarCraft introduced long-horizon goals, partial observability, and sparse rewards [kempka2016vizdoom, vinyals2017starcraft, fan2022minedojobuildingopenendedembodied, wang2023voyager]. With the rise of VLM/VLA agents, web-based benchmarks and browser environments have been proposed to test generalization to novel tasks and interfaces [yuan2023rl, qi2024webrl]. Following this line, we evaluate agents on a diverse suite of games and adopt simple game-agnostic metrics grouped into three levels—survival, curiosity, and utility—to provide their preformance from physical intuition to reasoning and their scaling with experience.
3 Preliminaries
Figure 3: IPR training pipeline. Stage 1: PhysCode pre-training. Video clips with optical flow and action semantics are fed to a VQ-based latent action model to learn discrete codes (PhysCode) that represent dynamics. Stage 2: Latent-conditioned world model. Given current features and PhysCode sequences, a world model is trained to predict future features and rewards under latent actions. Stage 3: Prediction-reinforced reasoning. A VLM reasons over the scene and generates candidate PhysCode sequences. The world model rolls them out in imagination, and the predicted rewards/values are used to select the best actions and to optimize the VLM policy.
3.1 Problem Setting
We consider a family of interactive environments $\{\mathcal{E}_{m}\}_{m=1}^{M}$, each formalized as a POMDP:
$\mathcal{M}_{m}=\big(\mathcal{S},\mathcal{A},T_{m},R_{m},\mathcal{O},\gamma;\,\varphi_{m}\big),$ (1)
where $\varphi_{m}$ are latent physics parameters (e.g., gravity $g$, friction $\mu$, mass $M$). At time $t$, the environment emits an image $x_{t}\!\sim\!\mathcal{O}(\cdot\!\mid\!s_{t})$, which we encode as $z_{t}=\phi_{\mathrm{enc}}(x_{t})$; the agent executes $a_{t}\!\in\!\mathcal{A}$ and transitions according to
$s_{t+1}\sim T_{m}\!\big(s_{t+1}\mid s_{t},a_{t};\varphi_{m}\big),\qquad r_{t}=R_{m}(s_{t},a_{t}),$ (2)
where physics resides in $T_{m}$, and causality in $R_{m}$.
Control may use one of several interfaces $A\in\{\textsc{keyboard},\textsc{language},\textsc{latent}\}$; a goal-conditioned VLM selects actions in the chosen space via
$a_{t}^{(A)}\sim\pi_{\omega}^{(A)}\!\big(\cdot\mid z_{t},\text{prompt}_{t}\big),\qquad a_{t}\equiv a_{t}^{(A)}\in\mathcal{A}.$ (3)
A feature-level world model $f_{\theta}$ then rolls out imagined futures under selected action sequences in the same action space $A$. Given a horizon $H\in\mathbb{N}$, initialize $\hat{z}_{t}:=z_{t}$ and choose an action sequence $\{a_{t+k}^{(A)}\}_{k=0}^{H-1}$. The rollout is defined by
$\hat{z}_{t+k+1}=f_{\theta}\!\big(\hat{z}_{t+k},\,a_{t+k}^{(A)}\big),\quad k=0,1,\ldots,H-1,$ (4)
where $k$ indexes the step inside the imagined trajectory from time $t$ to $t+H$.
3.2 PhysCode: Physics-centric Action Code
Motivated by the issues of raw-key semantic aliasing and the distortion of fine-grained visual dynamics when expressed in language, we propose PhysCode, a discrete latent action representation built on a VQ codebook $\mathcal{C}=\{v_{k}\}_{k=1}^{K}$. At step $t$, an action is a short code sequence $a_{t}^{\textsc{lat}}=\langle c_{t,1:L}\rangle$ with embedding obtained by looking up and pooling $\{v_{c_{t,\ell}}\}$.
Each code is conditioned on three cues: (i) domain-specific visual appearance via DINOv3 [siméoni2025dinov3] features $\phi_{\mathrm{img}}(x_{t})$, (ii) domain-agnostic motion via optical flow [fischer2015flownetlearningopticalflow] $\phi_{\mathrm{flow}}(\mathrm{Flow}(x_{t},x_{t+1}))$, and (iii) lightweight semantic hints extracted by a T5 encoder [raffel2023exploringlimitstransferlearning], with $\phi_{\mathrm{sem}}(y_{t})=\mathrm{Enc}_{\text{T5}}(y_{t})$. Since natural language alone cannot express fine-grained dynamics (e.g., impulse magnitude, frictional slip), we rely on flow and visual features to carry these details while keeping semantics as guidance. By design, the resulting codes capture physics-relevant intervention primitives that share across environments with similar underlying physics and separate when physics differ, enabling consistent reuse under matched physics and discrimination under shifted dynamics.
4 Method
In this section, we introduce three components of IPR (Fig. 3): (1) learning a physics-centric action code vocabulary across diverse physical principles and causal mechanisms; (2) training a latent-conditioned world model that predicts future features and rewards under sequences of latent actions; and (3) reinforcing VLM with world model rollout prediction in the interactive environment, using aligned latent action code. In inference, the VLM proposes candidate latent actions, queries the world model for short-horizon imagination and value estimates to score them, and executes the highest-scoring action.
Inducing the Latent Action Vocabulary.
Using the cues in Sec. 3.2 (DINOv3 appearance $f_{t},f_{t+\Delta}$, optical flow $u_{t}$, and lightweight semantics $e_{t}$), a small gated fusion module forms a fused representation $h_{t}$. A spatio-temporal encoder $E_{\psi}$ maps $h_{t}$ to a continuous code $z_{t}$, which is vector-quantized to an index $a_{t}\!\in\!\{1,\ldots,K\}$ with codebook $\mathcal{C}=\{c_{k}\}_{k=1}^{K}$, and a decoder $D_{\psi}$ predicts the future feature $\hat{f}_{t+\Delta}$ from $(f_{t},c_{a_{t}})$. We train with a standard VQ-VAE objective
$\mathcal{L}_{\mathrm{LA}}$ $=\big\|\hat{f}_{t+\Delta}-f_{t+\Delta}\big\|_{2}^{2}$ (5)
$\quad+\beta\big\|\mathrm{sg}[z_{t}]-c_{a_{t}}\big\|_{2}^{2}+\gamma\big\|z_{t}-\mathrm{sg}[c_{a_{t}}]\big\|_{2}^{2},$
augmented with modality dropout on flow and a mild gate-sparsity regularizer to avoid over-reliance on optional cues. Since optical flow is only available during pretraining, it acts as privileged information that helps shape a physics-centric codebook, while dropout and gate sparsity distill this structure into an encoder that, at test time, relies only on appearance and semantic cues. At inference, we disable the flow gate and reuse the same encoder to obtain $z_{t}$ and its quantized index $a_{t}$ from appearance+semantics only. The resulting discrete vocabulary yields temporally predictive tokens that cluster under matched physics and separate under different dynamics, providing a shared interface for VLM reasoning and world-model prediction.
Training the Latent-Level World Model with a Critic.
With the latent action vocabulary fixed, we train a feature-level world model to predict future features conditioned on latent actions, replacing raw controls with their PhysCode indices. For triples $(f_{t},a_{t},f_{t+\Delta})$, we embed $a_{t}$ to $e_{a_{t}}$ and compute
$(\hat{f}_{t+\Delta},\,V_{\theta}(f_{t},a_{t}))\;=\;P_{\theta}\big(f_{t},\,e_{a_{t}}\big).$ (6)
We predict in the latent space, since features compress appearance variance and rendering noise, making dynamics more shareable across games. Concretely, we first train the world model with a feature-prediction loss $\mathcal{L}_{\text{pred}}=\big\|\hat{f}_{t+\Delta}-f_{t+\Delta}\big\|_{1},$ and then learn a critic head with a Q-learning–style objective $\mathcal{L}_{\text{value}}=\ell_{\text{Q}}\!\big(V_{\theta}(f_{t},a_{t}),\,y_{t}\big),$ where $y_{t}$ is a target value computed from rollout returns via standard TD backups.
Prediction-Reinforced Interactive Reasoning.
We strengthen interactive reasoning with prediction: a world model imagines rollouts, and a VLM plans in the same latent action space. We adopt Qwen3-VL-8B [yang2025qwen3] as the backbone and extend its tokenizer with PhysCode tokens so the VLM can directly emit discrete latent actions while preserving its language ability.
We first align perception and action by supervised training on $(f_{t},c_{t})$ pairs, where $f_{t}$ is the DINOv3 feature of the current frame and $c_{t}$ the latent action learned in Stage 1. Given the current context and goal $g$, the VLM samples $B$ candidate PhysCode sequences $\{\mathbf{a}^{(b)}\}_{b=1}^{B}$, and the world model runs short-horizon imagined rollouts to assign each a predicted return, from which we compute advantages $A^{(b)}$. We then update the policy with GRPO [shao2024deepseekmath]:
$\mathcal{L}_{\text{GRPO}}=\frac{1}{B}\sum_{b=1}^{B}A^{(b)}\log\pi_{\phi}(\mathbf{a}^{(b)}\!\mid f_{t},g)\;-\;\beta\,\mathrm{KL}\!\big(\pi_{\phi}\,\|\,\pi_{0}\big),$ (7)
In inference, the VLM proposes latent action candidates, the world model scores and prunes them via short-horizon rollouts, and a router $T_{\text{env}}$ maps the selected PhysCode to environment controls. Through repeated interaction under this prediction-in-the-loop scheme, the experience collected from imagined and executed trajectories reinforces the VLM, improving its physical reasoning in interactive environments.
5 Experiments
Table 1: PhysCode validation. Left: Joint training across heterogeneous-physics games reveals cross-game conflicts for keyboard/mouse; language partially alleviates this via semantics, while PhysCode separates actions by dynamics, reducing interface aliasing and showing minimal degradation under physics shifts. Middle: Leave-$n$-out transfer: training on all but 10 titles and evaluating zero-shot on the held-out set, PhysCode transfers more reliably than keyboard or language interfaces. Right: Physics-conditioned transfer: zero-shot performance is relatively higher when target environments match the training set’s physical mechanisms, indicating that PhysCode captures reusable physical principles rather than game-specific bindings.
(a) Confusion test for joint training.
Latent-Predict Cosine $\uparrow$ MSE $\downarrow$ L1 $\downarrow$
Ad-hoc 0.9939 0.0121 0.0495
Keyboard 0.9894 0.0211 0.0772
Language 0.9892 0.0216 0.0758
PhysCode 0.9919 0.0204 0.0737
Pixel-Predict FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
Ad-hoc 87.83 0.7062 23.86
Keyboard 110.9 0.6110 20.82
Language 82.51 0.6960 23.52
PhysCode 80.35 0.7240 23.82
(b) Leave-$n$-out transfer.
Latent-Predict Cosine $\uparrow$ MSE $\downarrow$ L1 $\downarrow$
Pre-trained 0.9856 0.0230 0.0846
Keyboard 0.9784 0.0430 0.1153
Language 0.9790 0.0418 0.1132
PhysCode 0.9798 0.0403 0.1212
Pixel-Predict FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
Pre-trained 127.3 0.7438 22.11
Keyboard 315.0 0.3340 12.46
Language 320.2 0.1670 9.389
PhysCode 297.0 0.3533 13.04
(c) Physics-conditioned transfer.
In this section, we aim to answer three questions: (1) Why is PhysCode necessary compared with raw keyboard inputs or language instructions? (2) How would world model prediction reinforce VLM reasoning? (3) Would IPR show scaling potential to transfer to unseen games?
5.1 Setup: Datasets, Tasks, and Metrics
Sources.
We curate a multi-source benchmark covering 863 open-source retro titles (via stable-retro [stable-retro]), 134 lightweight HTML/Canvas games, and 3 commercial games. This breadth exposes agents to heterogeneous visuals, action interfaces, and underlying physics/causal mechanisms, encouraging models to capture shared physical–causal regularities rather than overfit to domain-specific biases.
Figure 4: Game data distribution. Our dataset spans over 1,000 games categorized by game category, control interface, operation and visual complexity, physical and causal mechanisms. This wide coverage enables agents to experience diverse domains and learn transferable physical and causal understanding.
Diversity axes.
We characterize each environment along seven axes to enable structured generalization analysis: (1) Game category, with emphasis on physical interaction (e.g., platformer, shooter, sports); (2) Control interface, such as GameBoy–style discrete keys, keyboard–mouse combinations, and high-dimensional hybrids; (3) Visual complexity, ranging from low-resolution pixel art to high-fidelity 3D; (4) View perspective, e.g. ego-centric, top-down, and side views; (5) Causal mechanism, e.g. damage/health dynamics, collection, punishment; (6) Physical principle, e.g. gravity, contact, and inertia; (7) Operational difficulty, approximated by the entropy and frequency of human control actions, reflecting how precisely and how often players must operate to succeed; Fig. 4 summarizes the distributions over sources, game types, and these axes; detailed per-environment statistics are provided in the supplementary.
Data collection and preprocessing.
Across the 1,000-game corpus, we record human play at 60 FPS for 4 minutes per title and obtain per-game annotations covering physical principles, causal mechanisms, action semantics, and game instructions. We perform a series of preprocessing, including normalizing time intervals, removing non-interactive segments, rebalancing extended idle/no-op periods, etc. More details are in the supplementary.
Hierarchical level design.
Inspired by Maslow’s hierarchy of needs [huitt2007maslow], we treat gameplay as a three-level progression: Survival $\rightarrow$ Curiosity $\rightarrow$ Utility (Fig. 2(a)), from intuition to reasoning.
[Uncaptioned image]Survival. The objective is to remain alive as long as possible, ignoring the original goal and avoiding risks. We report survival time normalized per game, $H=\mathbb{E}[T]/T_{\text{typ}}$, where $T$ is episode length (steps) and $T_{\text{typ}}$ is a per-game reference horizon (e.g., median survival under a random policy).
[Uncaptioned image]Curiosity. The goal is to visit novel states like a baby to uncover regularities in the environment’s dynamics and causal mechanisms. Following Magnipy [limbeck2024metric], we embed frames with a pretrained CLIP visual encoder [radford2021learningtransferablevisualmodels], compute the trajectory’s multi-scale metric-space magnitude curve $M(\tau)$, and define the exploration score as the area under this curve: $E=\mathrm{AUC}(M(\tau)),$ where larger $E$ indicates broader state-space coverage.
[Uncaptioned image]Utility. Utility measures how well an agent realizes Bentham’s utility of life [bentham1789morals]: devoting itself to goal completion with higher reward and shorter time. We evaluate downstream goals according to the game types (completion, score, checkpoint time) and report the human-normalized score (HNS) [bellemare2013ale] per game:
$\mathrm{HNS}=\frac{m-m_{\text{rnd}}}{m_{\text{hum}}-m_{\text{rnd}}}\,$, (8)
where $m$ is the agent metric, $m_{\text{rnd}}$ the random baseline, and $m_{\text{hum}}$ human performance.
Table 2: Comprehensive comparison across [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. Obj: training objective; Mean: normalized score; Avg. Rank: normalized average rank (1/(n1)1/(n{-}1), higher is better); Ratio@Top-3(%): percentage of games where the method is in the top-3.
Methods [Uncaptioned image]Survival [Uncaptioned image]Curiosity [Uncaptioned image]Utility
Mean Avg. Rank Ratio@Top-3(%) Mean Avg. Rank Ratio@Top-3(%) Mean Avg. Rank Ratio@Top-3(%)
Control Group
Random 0.541 0.450 3.5 5.254 0.363 5.8 0.000 0.559 1.1
Human 0.764 0.760 41.9 5.428 0.480 11.7 1.000 0.874 76.1
Imitation Learning (IL) Group
ACT-BC 0.541 0.474 5.8 6.896 0.476 13.0 0.092 0.487 5.7
Qwen3-VL-8B-BC 0.578 0.578 7.0 6.945 0.502 11.8 0.129 0.560 3.4
Reinforcement Learning (RL) Group
PPO@survival 0.571 0.581 16.3 5.453 0.415 2.3 0.259 0.636 14.8
PPO@curiosity 0.562 0.566 11.6 5.344 0.373 3.5 0.285 0.659 19.3
PPO@utility 0.554 0.609 17.4 5.419 0.397 2.3 0.268 0.637 18.2
DQN@survival 0.538 0.493 15.1 7.058 0.588 10.6 0.147 0.518 8.0
DQN@curiosity 0.562 0.531 18.6 6.841 0.578 12.8 0.159 0.521 8.4
DQN@utility 0.556 0.515 18.6 6.450 0.563 9.3 0.156 0.536 9.1
World Model Group
DreamerV3@survival 0.521 0.495 8.1 7.737 0.526 14.0 0.114 0.446 6.8
DreamerV3@curiosity 0.538 0.466 8.1 7.843 0.608 20.9 0.138 0.449 11.5
DreamerV3@utility 0.538 0.481 9.3 7.336 0.530 17.6 0.139 0.446 8.7
V-JEPA2@survival 0.531 0.394 5.8 7.815 0.469 14.0 0.065 0.285 4.0
V-JEPA2@curiosity 0.516 0.343 0.0 8.463 0.610 22.1 0.046 0.246 4.5
V-JEPA2@utility 0.526 0.369 1.2 7.777 0.412 15.1 0.051 0.293 4.5
GenieRedux@survival 0.539 0.435 3.5 7.937 0.573 17.4 0.090 0.334 4.0
GenieRedux@curiosity 0.534 0.421 3.5 8.390 0.675 20.9 0.084 0.293 3.4
GenieRedux@utility 0.540 0.445 4.7 8.067 0.573 17.4 0.085 0.349 4.5
Multimodal Large Language Model (MLLM) Group
GPT-4o@survival 0.575 0.605 14.0 4.954 0.390 2.2 0.156 0.543 7.0
GPT-4o@curiosity 0.521 0.429 7.0 5.800 0.518 4.7 0.124 0.474 4.5
GPT-4o@utility 0.522 0.474 8.1 5.676 0.465 2.3 0.162 0.549 8.0
GPT-5@survival 0.619 0.630 25.6 5.181 0.358 2.3 0.199 0.593 18.2
GPT-5@curiosity 0.529 0.488 5.8 5.621 0.448 5.8 0.148 0.515 5.7
GPT-5@utility 0.536 0.460 5.8 5.329 0.427 2.3 0.206 0.606 13.6
Qwen3-VL-A30B@survival 0.544 0.514 5.8 5.690 0.522 5.8 0.149 0.539 6.8
Qwen3-VL-A30B@curiosity 0.518 0.450 7.0 7.113 0.600 11.7 0.133 0.500 4.5
Qwen3-VL-A30B@utility 0.545 0.504 9.3 6.214 0.443 4.7 0.146 0.552 6.8
Interactive Physical Reasoner
IPR (8B) 0.589 0.517 10.4 7.874 0.584 15.6 0.178 0.604 8.9
(ranking) (2/26) (8/26) (9/26) (5/26) (6/26) (7/26) (7/26) (5/26) (9/26)
Key Takeaways across [Uncaptioned image]Survival, [Uncaptioned image]Curiosity, and [Uncaptioned image]Utility Prediction-based Methods (WM). Strong at [Uncaptioned image], but weaker at [Uncaptioned image] and [Uncaptioned image]. Trained on broad exploratory trajectories, latent rollouts broaden coverage and reveal dynamics, but tend to imitate visually-alike futures rather than reliably pursue goals. So prediction is useful as a look-ahead prior for risk and candidate actions. RL-based Methods (PPO, DQN). Strong at [Uncaptioned image]and [Uncaptioned image] when rewards are well-shaped, but weaker on [Uncaptioned image] and tasks without explicit goals. Reward gradients enable effective credit assignment under the right signal, yet sparsity and partial observability induce instability and interface overfitting—so RL works best as an optimization method. Experience-based Methods (Behavior Cloning). Strong at human-like [Uncaptioned image], but weaker on [Uncaptioned image] and [Uncaptioned image]. Deliberately imitate human trajectories and thus excel at low-risk survival, but struggle once tasks require precise control or exploration, and their performance depends strongly on the coverage and quality of the demonstrations. Reasoning-based Pretrained VLMs. Strong at goal-conditioned [Uncaptioned image]and [Uncaptioned image]; weaker on [Uncaptioned image]. They excel at instruction-driven reasoning but cannot predict consequences in the visual state space, so they work best as high-level reasoners that need auxiliary prediction modules for outcome-aware decisions. Interactive Physical Reasoning (Ours). Robust across [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. We combine the strengths of all three paradigms: VLMs provide goal-driven causal reasoning, the world model supplies rollout prediction, and RL optimizes decisions using imagined rewards, yielding consistently strong performance across all three levels.
5.2 Why is PhysCode Necessary
We first investigate whether PhysCode is necessary compared with raw keyboard/mouse inputs and natural-language instructions. First, we assess robustness under mixed-game joint training with heterogeneous physics (Tab. 1(a)), examining which action space best performs in diverse physical mechanisms and different console/game interfaces. Second, we test transfer (Tab. 1(b), Tab. 1(c)): a shared PhysCode learned on source games improves zero-shot performance in unseen environments with matched physics, demonstrating genuine physics grounding rather than interface memorization.
First, we examine how different action spaces behave when trained jointly across a mixture of games with heterogeneous physics (Tab. 1(a)). In this regime, raw keyboard/mouse inputs exhibit cross-game conflicts (the same key triggers different behaviors across environments). Language interfaces partially alleviate this via explicit semantics. PhysCode separates actions by dynamics, reducing interface aliasing and showing minimal degradation under physics shifts.
Next, we ask whether sharing the latent space supports transfer. In a leave-$n$-out protocol (Tab. 1(b)), we train on all but 10 games and evaluate zero-shot on the held-out titles. We find that PhysCode transfers more reliably than keyboard or language instructions.
Moreover, we condition transfer on the physics of the environment. We group games by their dominant physical mechanism, train under one principle (e.g., gravity), and evaluate zero-shot on held-out games with matching or different mechanisms. When targets match the training physics, zero-shot performance is typically higher (Tab. 1(c)), with notable exceptions such as inertia, which may already be covered by projectile/impulse. This suggests that PhysCode captures reusable physical mechanisms rather than game-specific bindings, even though our coarse physics taxonomy does not perfectly align with the agent’s internal abstractions.
5.3 Playing in Diverse Physical Worlds of Games
We evaluate IPR against prevalent baselines on 200 games, chosen to match the full dataset’s distribution of types, action spaces, and physics/causality. The baselines include:
  • RL. We utilize Multitask PPO [yu2020meta] (policy-based) and shared-parameter DQN [osband2016deep] (value-based) as standard reinforcement learning approaches.
  • VLM. We employ a range of vision-language models, including closed-source models such as GPT-4o and GPT-5 [openai2024gpt4ocard], as well as open-source models like Qwen3-VL-30B-A3B [yang2025qwen3].
  • World Model. We compare three different world models: DreamerV3 [hafner2023mastering] (latent-based), V-JEPA2 [assran2025v] (pretrained latent-based prediction), and Genie [bruce2024genie] (pixel-based prediction) (we follow GenieRedux implementation [kazemi2024learninggenerativeinteractiveenvironments]).
  • IL. We apply imitation learning (IL) models, including ACT [zhao2023learning] (end-to-end model) and Qwen3-VL-8B [yang2025qwen3] (VLM-based model).
We assess every model on the three hierarchical objectives, instantiating level-specific training or prompting. Further implementation details are provided in the supplementary. The key results are reported in Tab. 2. Takeaways are below the table.
5.4 Zero-shot Transferring to Unseen Games
To validate our Games-to-Unseen (G2U) setting, we construct a held-out target set $\mathcal{T}_{\mathrm{U}}$ of 50 games that are never used for training. From the remaining pool, we form stratified training subsets $\{\mathcal{S}_{N}\}$ of increasing size $N$, balanced by physics and causal mechanisms to control for domain bias. For each $N$, we train our IPR paradigm end-to-end on $\mathcal{S}_{N}$ and directly evaluate zero-shot on $\mathcal{T}_{\mathrm{U}}$ without any adaptation or reward re-scaling.
Across all three objectives, performance increases steadily with $N$, with the steepest early gains on [Uncaptioned image], followed by sustained improvements on [Uncaptioned image] and [Uncaptioned image] as more diverse interactions are observed. This suggests that training in physically and causally related environments helps IPR move beyond domain-specific quirks (visual style, control interface) and focus on shared physical and causal patterns (e.g., gravity, contact, momentum). In other words, as interactive experience accumulates, IPR behaves more human-like: it carries over physical priors and causal expectations rather than memorizing domain appearance or controls, demonstrating potential to further scale in richer interactive domains.
Figure 5: G2U zero-shot scaling on 50 held-out games. As the number of training games NN increases, zero-shot performance on Refer to caption, Refer to caption, and Refer to caption improves steadily on the unseen set $\mathcal{T}_{\mathrm{U}}$.
5.5 Ablations and Analysis
Does prediction help VLM reasoning?
[…]
Table 3: Ablation study results for IPR components of World Model prediction and GRPO.
Method [Uncaptioned image]Survival [Uncaptioned image]Curiosity [Uncaptioned image]Utility
VLM (pretrained) 0.62 2.14 0.89
VLM + BC 0.63 1.88 0.87
VLM + PPO 1.00 1.79 1.23
VLM + BC + PPO 0.57 1.86 0.77
IPR 0.76 2.77 1.34
6 Discussion
We study an interactive physical reasoning paradigm in which a general-purpose VLM reasons in language, acts through a physics-centric latent interface (PhysCode), and is reinforced by imagined rewards from a world model, asking whether such agents can internalize physical and causal regularities from heterogeneous games and show clear scaling as experience grows. From this perspective, latent-action world models (e.g. Genie, UniVLA [bruce2024genie, bu2025univlalearningacttaskcentric]) learn discrete action abstractions and latent dynamics for controllable rollouts; imagination-based control methods (e.g. Dreamer, V-JEPA2-AC [hafner2024masteringdiversedomainsworld, assran2025vjepa2selfsupervisedvideo]) optimize policies inside learned world models over device-level actions; and large-scale VLM-based game agents (e.g. Game-TARS [wang2025gametarspretrainedfoundationmodels]) scale vision–language–action models with massive human demonstrations and auxiliary multimodal tasks. Yet, from a physics-centric perspective, these approaches do not explicitly organize actions by shared physical mechanisms across hundreds of games or align VLM’s reasoning ability with prediction competence in a common latent space. IPR combines their advantages to study how physical knowledge and transfer emerge under the unified Survival-Curiosity-Utility evaluation, though it is still limited to game environments and short-horizon imagination, leaving real-world transfer and longer-horizon reasoning to future work.
7 Conclusion
[…]